I am really confused on the following problems.
Find the equation, in standard form, for the line that has the following characteristics.
d. (5, 2); m = -1/2
e. (2,-1); (-1, 3)
for letter e. how do you use the formula m= x2-x1/ y2-y2?
any way of explaining the problems would be a big help!

Answer by troyapplehelen(46)

1. the line formular is y=mx+b
in d, m is -1/2, y=(-1/2)x+b
passing the point (5,2), 2=(-1/2)5+b
=(-5/2)+b then b=9/2
answer) y-(-1/2)x+(9/2)
2. the formular for the slope is (change of y)/(change of x)
(u got the wrong formular).
so, m is 4/(-3), then y=(-4/3)x+b
to find value of "b", substitude any point value.
3=(-4/3)(-1)+b
=(4/3)+b, then, b is 5/3
answer) y=(-4/3)x+(5/3)